OpenStudy (anonymous):

Write an equation in slope intercept form fro the line described. Slope -6/5, passes through (0,8)

OpenStudy (anonymous):

well your basic equation would be y=(-6/5)x+8

OpenStudy (anonymous):

so would m=-6/5, b=8 y=-6/5+8 ?

OpenStudy (anonymous):

Yes, but do not forget to add an 'x' to the last equation. You just put y=-6/5 + 8. The 'x' added to the -6/5 signifies the slope.

OpenStudy (anonymous):

Ok thanks

OpenStudy (anonymous):

You're welcome!

OpenStudy (anonymous):

how about slope -2, passes thru (-3,14) would it be y=2x+8 ?

OpenStudy (anonymous):

I'm not quite sure where you are getting the 8 from. You're slope is -2 so it should be y= -2x instead of just 2x. Be careful with your signs. The b in the equation should just be 14. Ultimately, the equation should be y=-2x+14